OpenStudy (anonymous):

The number of men in a tour group is 15 times the number of women. If there are 48 people in the group, how many are men?

Directrix (directrix):

Let M = the number of men on the tour Then, let W = the number of women on the tour M + W= 48 But, M = 15 * W because the number of men is given to be 15 times the number of women. Substitute 15W for M in the equation: M + W = 48. Solve for M. Then, back substitute what you get into M = 15 * (W). Post what you get, okay? @zxcvbnm08765

Directrix (directrix):

15W + 1W = 48 What is the value of W?

OpenStudy (anonymous):

would it be 30 @Directrix

Directrix (directrix):

No. 15W + 1W = 48 16W = 48 W = 3 So, 3 women. Multiply 3 by 15 to get the number of men.

OpenStudy (anonymous):

oh ok 45

Directrix (directrix):

That is what I got.